OpenStudy (anonymous):

What is sin Pi/4 in a unit circle?

OpenStudy (anonymous):

please help!

jimthompson5910 (jim_thompson5910):

use this unit circle http://www.regentsprep.org/Regents/math/algtrig/ATT5/600px-Unit_circle_angles_svg.jpg locate the angle pi/4 then find the corresponding point on the unit circle for angle pi/4 what is the y coordinate of that point?

OpenStudy (anonymous):

is it 45 degrees?

jimthompson5910 (jim_thompson5910):

45 degrees = pi/4 radians the final answer isn't 45 degrees though

OpenStudy (anonymous):

i'm not sure what to do after then....

jimthompson5910 (jim_thompson5910):

look at the attached image notice how I'm circling the point when the angle is pi/4

jimthompson5910 (jim_thompson5910):

what point did I circle? what is the y coordinate of that point?

OpenStudy (anonymous):

sqrt 2/2?

jimthompson5910 (jim_thompson5910):

therefore, \[\Large \sin\left(\frac{\pi}{4}\right) = \frac{\sqrt{2}}{2}\]

jimthompson5910 (jim_thompson5910):

the y coordinate corresponds to sine the x coordinate corresponds to cosine

OpenStudy (anonymous):

so the answer is Pi/4?

OpenStudy (anonymous):

the answer is 1/root2

OpenStudy (anonymous):

sorry, i meant sqrt 2

jimthompson5910 (jim_thompson5910):

Another example, \[\Large \sin\left(\frac{\pi}{3}\right) = \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

i understood, i just typed the wrong thing that iwas trying to say. but thank you!

jimthompson5910 (jim_thompson5910):

no problem